OpenStudy (anonymous):

Factor completely. 6a^2 + 19as + 15s^2

OpenStudy (chaise):

(2a+3s)(3a+5s) Do you need help with how this answer was obtained?

OpenStudy (anonymous):

Yes

OpenStudy (chaise):

It's a little bit of trial and error, and very hard to explain. After a while you notice relationships between numbers.

OpenStudy (anonymous):

Yes, that was discussed in class. Isn't this called the FOIL technique or something like that?

OpenStudy (anonymous):

If you multiply the lead coefficient, 6, time the constant, 15, you get 6*15=90. The poly will be factorable if there are factors of 90 that add up to 19, the coefficient of the middle term. In this case: 9*10=90 9+10=19 So you can use the 9 and the 10 to rewrite the 19as to be 9as+10as.\[6a^2+9as+10as+15s^2\]This can be factored by"grouping method" in which you factor out the GCF of the first two terms and the GCF of the second two terms:\[3a(2a+3s)+5s(2a+3s)\]Now there are two terms in which the binomial 2a+3s is a GCF, so factor it out to get\[(2a+3s)(3a+5s)\]And you are done.